OpenStudy (koolkid98):

Jenna buys 12 lb of meat. Some of the meat is chicken and costs $1.90/lb. The rest is beef, which costs $4.30/lb. She spends a total of $34.80. Which equation models this situation? Let c represent the number of pounds of chicken she buys. A. (1.9 + 4.3)(12 – c) = 34.8 B. 4.3c + 1.9(c + 14) = 34.8 C. 4.3c + 1.9(12) = 34.8 D. 1.9c + 4.3(12 − c) = 34.8

OpenStudy (koolkid98):

@gnech

OpenStudy (koolkid98):

i wanna say c

OpenStudy (danjs):

You can use the information to make two equations, one for the pounds, and one for the cost. Total Pounds = Chicken + Beef 12 = C + B Total Cost = Chicken Lbs * cost per lb + Beef Lbs * Cost per lb 34.80 = 1.90*C + 4.30*B

OpenStudy (koolkid98):

what is it

OpenStudy (danjs):

12 = C + B 34.80 = 1.90*C + 4.30*B Here they went a step further and solved the first one for B, and substituted that B into the second one... B = 12 - C so 34.80 = 1.90*C + 4.30*(12 - C)

OpenStudy (koolkid98):

what?

OpenStudy (koolkid98):

@DanJS

OpenStudy (danjs):

That is how you do the problem, the last line is what they want...

OpenStudy (koolkid98):

ok